An alpha particle, with velocity of magnitude 500 m/s, travels horizontally into a region of space in which

there is an upward-directed uniform electric field of magnitude 29 N/C and a uniform magnetic field. The

magnetic field makes an angle of 35° with the particle’s velocity vector and is such that it causes the alpha

particle to pass through the region of space undeflected.


How would you determine the magnitude of the magnetic field that is required to cause the alpha particle to pass

through undeflected?



1
Expert's answer
2021-04-19T07:24:22-0400

To be given in question

Vx=500 meter/sec

Electric field==29 N/C

"\theta" =35°

To asked in question

Magnetic field =?



"F=qE \longrightarrow (1)"

"F=q(V\times B) \longrightarrow(2)"


equation (1)and(2)ate equal

"qE =q(V\times B)"

"E=VBsin\theta"

Put value

"B=\frac{29}{500}sin35°"

"B=0.03326 Tesla"